OpenStudy (anonymous):

Compress into one logarithm. 3log(x+1) - 4[log x + log(x^2-1)]

OpenStudy (anonymous):

Well your first step is to use one of the properties of log which states that a number before log and be raised to a power. So in this problem you could take 3 from log(x+1) and raise it as a power so it becomes log(x+1)^3

OpenStudy (anonymous):

There are two other properties you can use in this problem. \[\log(M)- \log(N) = \log(m/N)\] and another \[\log(M) + \log(N) = \log(MN) \] Hope this helps